OpenStudy (anonymous):

a gun has a muzzle speed of 100 m/s. what angle of elevation should be used to hit an object 200 m away? assume the gun was fired from ground level. neglect air resistance and use g = 9.8 m/s^2 as the acceleration of gravity

OpenStudy (anonymous):

assume:a=angle h=height t=time x=distance since h=0.5g*t^2 and the distance x=v*cos(a)*t=100cos(a)*t then we can build a parabola function which when h=0, x1-x2=200

OpenStudy (anonymous):



OpenStudy (anonymous):

almost forgot you have to caculate the initiative vertical velocity which is 100sin(a)